| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  164  
 | 
13 Sep 2024, 02:16 | 
Wonhyuk Jo | MID | Shift summary | - DAQ was crashed after taking darks of 
AGIPD. It was solved by DOC and OCD.
 
 
- Laser could not switched from | 
   |  | 
  163  
 | 
12 Sep 2024, 23:48 | 
Adam Round | SPB | Shift end | X-ray delivery
 
 
8.951 
keV, 1.6 mJ | 
   |  | 
  162  
 | 
10 Sep 2024, 07:26 | 
Romain Letrun | SPB | Shift summary | X-ray 
delivery
 
 
 
	12.8 keV, 1-20 pulses at | 
   |  | 
  161  
 | 
09 Sep 2024, 23:19 | 
Peter Zalden | FXE | Issue | Finally, the drift velocity of the pointing 
reduced approx a factor of 2 as compared 
to yesterday, see attached. Possibly it will | 
    |  | 
  160  
 | 
09 Sep 2024, 06:52 | 
Romain Letrun | SPB | Shift summary | X-ray 
delivery
 
 
 
	12.8 | 
   |  | 
  159  
 | 
08 Sep 2024, 20:30 | 
Peter Zalden | FXE | Issue | We observe a continuous horizontal beam 
drift in SA1 after the M1 incident. It seems 
not to converge, so one can wonder where | 
    |  | 
  158  
 | 
08 Sep 2024, 06:44 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
12.8 keV, ~0.8 mJ
 
 
Achievements/Observations | 
   |  | 
  157  
 | 
30 Aug 2024, 13:02 | 
Antje Trapp | XRO | M1 Rz damaged - DO NOT MOVE!!! | What has happened:
 
During euler motion calibration of 
M1, we were driving the vertical motors by | 
 6x   |  | 
  156  
 | 
27 Aug 2024, 12:17 | 
Chan Kim | SPB | Shift summary | X-ray delivery
 
 
 
	12.78 keV, ~1.3 mJ
 
 
 
Optical 
laser delivery | 
   |  | 
  155  
 | 
26 Aug 2024, 23:05 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
-12.40 - 12.82 keV, ~1.3 mJ
 
 
Optical laser delivery | 
   |  | 
  154  
 | 
25 Aug 2024, 22:37 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
-12.40 
- 12.82 keV, XDT9 XGM ~0.85 mJ, 3 hours interruption | 
   |  | 
  153  
 | 
24 Aug 2024, 23:09 | 
Chan Kim | SPB | Shift summary | X-ray delivery
 
 
-12.40 
- 12.82 keV, XDT9 XGM ~0.81 mJ@ end of shift | 
   |  | 
  152  
 | 
23 Aug 2024, 08:48 | 
Peter Zalden | FXE | Shift summary | Used 9.4 keV with 94 kHz wed and thu nights. 
Received approx. 0.9 mJ with large fluctuations 
in pulse energy, however acceptable for the | 
   |  | 
  151  
 | 
22 Aug 2024, 23:15 | 
Raphael de Wijn | SPB | Shift summary | X-ray delivery
 
 
12.40 - 12.82 keV, ~1.1 mJ
 
 
Optical laser delivery | 
   |  | 
  150  
 | 
21 Aug 2024, 23:34 | 
Raphael de Wijn | SPB | Shift summary | X-ray delivery
 
 
12.40 - 12.82 keV, ~1.1 mJ
 
 
Optical laser delivery | 
   |  | 
  149  
 | 
13 Aug 2024, 02:04 | 
Wonhyuk Jo | MID | Shift summary | Found the interference fringe at Suna microscope 
with Zyla cam from MID split-and-delay line. 
 
The fringe patterns were recorded | 
   |  | 
  148  
 | 
12 Aug 2024, 23:47 | 
Raphael de Wijn | SPB | Shift summary | 
 
X-ray 
delivery
 
 
11.56 keV, 202 pulses @0.5 MHz, | 
   |  | 
  147  
 | 
12 Aug 2024, 02:07 | 
Wonhyuk Jo | MID | Shift summary | Found spatial overlap at Suna microscope 
with Zyla cam.
 
The splitting ratio was obtained about | 
   |  | 
  146  
 | 
11 Aug 2024, 23:07 | 
Raphael de Wijn | SPB | Shift summary | X-ray delivery
 
 
11.56 keV 202 pulses @0.5 MHz, 
~1.0 mJ | 
   |  | 
  145  
 | 
11 Aug 2024, 14:33 | 
Romain Letrun | SPB | Shift summary | Summary 06.08 
- 10.08
 
 
X-ray | 
   |   
 |